# Philip ${}^{*\,\sharp}$ Yandex, MIPT &Aleksandr Beznosikov

### v1

@@ Abstract
With the increase in the number of parameters in large language models, the process of pre-training and fine-tuning increasingly demands larger volumes of GPU memory. A significant portion of this memory is typically consumed by the optimizer state. To overcome this challenge, recent approaches such as low-rank adaptation (LoRA (Hu et al., 2021)), low-rank gradient projection (GaLore (Zhao et al., 2024a)), and blockwise optimization (BAdam (Luo et al., 2024)) have been proposed. However, in all these algorithms, the effective rank of the weight updates remains low-rank, which can lead to a substantial loss of information from the gradient. This loss can be critically important, especially during the pre-training stage. In this paper, we introduce FRUGAL (Full-Rank Updates with GrAdient spLitting), a new memory-efficient optimization framework. FRUGAL leverages gradient splitting to perform low-dimensional updates using advanced algorithms (such as Adam), while updates along the remaining directions are executed via state-free methods like SGD or signSGD (Bernstein et al., 2018). Our framework can be integrated with various low-rank update selection techniques, including GaLore and BAdam. We provide theoretical convergence guarantees for our framework when using SGDM for low-dimensional updates and SGD for state-free updates. Additionally, our method consistently outperforms concurrent approaches across various fixed memory budgets, achieving state-of-the-art results in pre-training and fine-tuning tasks while balancing memory efficiency and performance metrics.
\$*\$\$*\$footnotetext: Work was mostly done during the internship at Mohamed bin Zayed University of Artificial Intelligence.$\sharp$\$$\sharp$\$footnotetext: The corresponding author: zmushko.ph.a@gmail.com\$$\diamond$\$$\diamond$\$footnotetext: Moscow Institute of Physics and Technology §$\S$ Mohamed bin Zayed University of Artificial Intelligence\$$\dagger$\$$\dagger$\$footnotetext: Ivannikov Institute for System Programming of the Russian Academy of Sciences $\ddagger$ Innopolis University
@@ 1 Introduction
In recent years, Large Language Models (LLMs) such as GPT (OpenAI, 2023) and LLaMA-3 Dubey et al. (2024) have demonstrated remarkable performance across various disciplines (Brown, 2020; Yang et al., 2024; Romera-Paredes et al., 2024). However, a critical factor in achieving these results is the size of these models (Hoffmann et al., 2022). A larger number of parameters not only increases computational cost but also significantly raises memory requirements. For instance, training an 8 billion parameter LLaMA model in a 16-bit format necessitates each parameter to occupy 2 bytes, resulting in 16GB for storing the parameters and an additional 16GB for gradients. Utilizing the Adam optimizer (Kingma, 2014), which is standard for pre-training and fine-tuning LLMs, adds a further 32GB of memory to store the $m$ and $v$ statistics, resulting in 64GB total amount of memory. Furthermore, to achieve higher-quality results, training in pure 16-bit format is often insufficient (Zamirai et al., 2020). This necessitates storing master weights and optimizer statistics in 32-bit format, leading to total memory demands that exceed the capacity of cutting-edge graphics cards, such as the A100-80GB.
Numerous research projects have been aimed at reducing these significant costs. These approaches include engineering solutions like gradient checkpointing Chen et al. (2016) and memory offloading (Rajbhandari et al., 2020), which do not change the training trajectory. There are also methods that adjust the training algorithm by decreasing the number of trainable parameters (Frankle & Carbin, 2018; Wang et al., 2023; Sreenivasan et al., 2022; Horváth et al., 2024) or their bit precision (Wortsman et al., 2023), as well as optimizer statistics (Dettmers et al., 2021; Shazeer & Stern, 2018; Zhang et al., 2024c). In this work, we concentrate on the latter category.
Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA (Hu et al., 2021), Dora (Liu et al., 2024), and BitFit (Zaken et al., 2021) reduce memory costs by training a relatively small number of parameters compared to the size of the original model, while the remaining modules are frozen. This approach has proven effective for the task of efficient fine-tuning of pre-trained language models. However, PEFT methods have a fundamental limitation: parameter updates always lie in a low-dimensional subspace, which prevents the use of these methods for pre-training (Lialin et al., 2023) and may constrain their capabilities in fine-tuning (Zhang et al., 2024a).
Recent works, such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023), and BAdam (Luo et al., 2024), offer a solution to this problem. These methods enable higher-dimensional full-parameter learning by periodically changing the optimizable low-rank subspace, which we will call the state-full subspace. However, even though these methods result in overall parameter changes that are high-dimensional, the updates in each step remain low-dimensional. The dimensionality of the state-free subspace significantly exceeds that of the state-full subspace. The remaining information contained in the gradient is not utilized for parameter updates. Nevertheless, this information can still be leveraged to train the model.
@@
Input: model $f_{\mathbf{\theta}}(\cdot)$ with $p$ parameters $\{\theta_{i}\in\mathbb{R}^{d_{i}}\}_{i=1}^{p}$, loss function $\mathcal{L}$, gradient projectors $P_{k,i}$ for $i\in[p],k\in[K]$.
@@
3:     compute $\ell\leftarrow\mathcal{L}(f_{\theta}(x),y)$triangleright Forward
@@
5:         $g_{\text{full},i}\leftarrow P_{k,i}(g_{i}),\;$triangleright Project Grad
6:         $g_{\text{free, i}}\leftarrow g_{i}-P^{-1}_{k,i}(g_{\text{full},i})$triangleright Residual
7:         $\texttt{s}_{\theta_{i}}\leftarrow[P_{k,i}(P^{-1}_{k-1,i}(s),\;s\in\texttt{s}_{\theta_{i}}]$$\triangleright$ Project States
@@
We present the FRUGAL framework, designed to bridge this gap. Our approach stems from a crucial observation: although memory constraints prevent the use of advanced optimizers like Adam (Kingma, 2014) in the state-free subspace, one still can update this subspace using state-free optimization algorithms like Stochastic Gradient Descent (SGD) or signSGD (Bernstein et al., 2018). This solution allows for high-dimensional updates, which provides additional opportunities to explore the parameter space and improves convergence.
  • We provide theoretical convergence guarantees for our framework. In the proof, we consider the case where SGDM acts as the state-full optimizer and SGD as the state-free optimizer, and we show that FRUGAL matches the best-known convergence rate in many scenarios.
  • To verify the practical applicability of FRUGAL, we conduct extensive experiments in popular real-world scenarios. [^1] The code is available at https://github.com/fzmushko/FRUGAL. In these experiments, we pre-train LLaMA-like models (up to 1B parameters) on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020) and fine-tune RoBERTa (Liu, 2019) on the GLUE benchmark (Wang, 2018). The results show that our method significantly outperforms previous memory-efficient algorithms while using the same memory budget.
  • We demonstrate that only the Logits layer in transformer-like models requires advanced optimizers like Adam, while other modules (including Embeddings and RMSNorms) can use simpler methods like signSGD without significant performance loss. This opens new possibilities for memory-efficient training and provides crucial insights into the learning dynamics of Transformers.
@@ 2 Related work
Memory-efficient full-parameter learning. Recent research has focused on reducing the memory footprint of LLMs by decreasing the size of optimizer states while maintaining their performance. Low-rank adaptation methods, such as LoRA (Hu et al., 2021), inject trainable rank decomposition matrices into each layer of the model, reducing memory requirements by optimizing only a few learnable adapters. ReLora (Lialin et al., 2023) builds upon this by merging low-rank adaptations into the main model weights during training, potentially increasing the total rank of the update. BAdam (Luo et al., 2024) leverages Block Coordinate Descent for full-parameter training by switching active blocks during fine-tuning. MicroAdam (Modoranu et al., 2024) compresses gradient information before feeding it into the optimizer state, significantly reducing memory footprint while enabling full parameter learning through error feedback mechanisms. GaLore (Zhao et al., 2024a) maintains full parameter learning by projecting gradients onto a low-rank subspace using truncated SVD decomposition, storing optimizer states in this reduced space. Notably, GaLore achieves good results in pre-training, with performance close to that of Adam. However, while these methods effectively reduce memory overhead, they all perform low-rank updates at each iteration. In contrast, our approach utilizes all available gradient information to perform full-dimensional updates at each optimizer step, offering a novel perspective on memory-efficient optimization for LLMs.
@@
Other memory-efficient optimization. Several other methods have been proposed to reduce the memory footprint of optimizers. AdaFactor (Shazeer & Stern, 2018) attempts to mimic Adam’s behavior while reducing memory usage through factorization of the variance matrix $v$. Adam-mini (Zhang et al., 2024c) further reduces memory by storing only one value $v$ per block. Dettmers et al. (2021) and Li et al. (2024) decrease memory footprint by quantizing optimizer states to the lower-precision representations. Lv et al. (2023) proposed to reduce weight gradient memory by fusing the backward operation with the optimizer update. Notably, these approaches are orthogonal to our method FRUGAL and can be combined with it for further memory efficiency.
@@
Block Coordinate Descent. Block Coordinate Descent (BCD) is a well-established optimization method with a rich history in mathematical optimization (Ortega & Rheinboldt, 2000; Tseng, 2001; Richtárik & Takáč, 2014; 2015b; Richtárik & Takác, 2016; Takáč et al., 2013; Richtárik & Takáč, 2015a). In recent years, a specific instance of BCD, known as layer-wise learning, has been applied to deep learning. Notable examples include (Luo et al., 2024; Pan et al., 2024), which leverage this approach for LLM fine-tuning. To the best of our knowledge, our work presents the first theoretical analysis of an extended BCD framework (Section 4) where the remaining coordinates are also updated using a different algorithm. This novel approach extends traditional BCD techniques, opening new avenues for full model optimization in deep learning.
@@
Sign-based methods for training language models. Since its introduction, Adam has become the de facto primary optimization algorithm, demonstrating superior practical results compared to SGD-based algorithms across various deep learning tasks. This difference is particularly noticeable when training Transformers on language tasks. While Zhang et al. (2020) hypothesized that Adam outperforms SGD in this setup due to the heavy-tailed distribution of sampling-induced errors, Kunstner et al. (2023) demonstrated that this superiority persists even in full-batch training. They proposed a new hypothesis suggesting that Adam’s key success factor is related to its similarity to signSGD (Balles & Hennig, 2018; Balles et al., 2020), and both Kunstner et al. (2023) and Zhao et al. (2024b) showed that signed descent with momentum reduces the performance gap with Adam. In contrast, to the best of our knowledge, we are the first to train the majority of language model’s parameters using signSGD without momentum, achieving minimal loss in quality. This approach further demonstrates the effectiveness of sign-based methods for LLM training, paving the way for more efficient and scalable optimization strategies in deep learning.
Table 1: Comparison of different projection and optimization strategies on pre-training LLaMA-130M on C4 with Adam as the state-full algorithm. State-free subspace is updated for full-rank methods but frozen for low-rank methods. Low-rank SVD outperforms Low-rank Random at the beginning of training but falls behind later due to insufficient exploration of the parameter space.
Method Validation perplexity after iterations $\downarrow$
4k 20k 40k 100k 200k
Low-rank SVD 39.75 26.35 24.38 22.30 21.11
Low-rank Random 42.31 25.99 23.55 21.33 20.01
Full-rank Random 37.26 23.46 21.53 19.66 18.64
Full-rank SVD 33.96 22.54 21.01 19.30 18.35
Full-rank RandK 36.38 23.02 21.25 19.70 18.63
Full-rank Block 37.20 23.34 21.42 19.59 18.60
Adam 33.95 21.90 20.56 18.97 18.13
3 FRUGAL: Full-Rank Updates with GrAdient spLitting
In this section, we first introduce two main building blocks of our method and then introduce our general framework.
@@ 3.1 The importance of exploring the entire space during the training process
In recent work, Zhao et al. (2024a) proposed GaLore, an optimization method based on projecting the gradient matrix ${\bm{G}}$ of each layer onto a low-dimensional subspace. To obtain the projection matrix ${\bm{P}}$, they use SVD decomposition of ${\bm{G}}_{t}$, which is recomputed with frequency $T$. The vectors or rows of ${\bm{G}}$ are projected onto the first $r$ left or right singular vectors, respectively. This approach has theoretical foundations: the first $r$ singular vectors correspond to the first $r$ singular values and, therefore, should better utilize information from the spectrum of ${\bm{G}}$.
@@
The authors pointed out that calculating the SVD decomposition results in extra computational overhead, which can be as much as a 10% increase as the hidden size of the model grows. To minimize this cost and examine the significance of using SVD decomposition, one may wonder about the possibility of employing a random semi-orthogonal projection matrix ${\bm{R}}$ instead of projecting onto the first $r$ singular columns with ${\bm{P}}$. Surprisingly, we found that although SVD decomposition delivers an initial boost, subsequent training with random projection yields significant improvements. As an illustration, we took the pre-training [^2] See Section 5.1 for a detailed description and discussion on the experimental setup. of a 130M model with LLaMA-like architecture on the C4 dataset (Raffel et al., 2020). The results are presented in Table 1, where training with ${\bm{P}}$ corresponds to the row labeled “Low-rank SVD”, and training with ${\bm{R}}$ corresponds to the row labeled “Low-rank Random”. The ranks of both projections ${\bm{P}}$ and ${\bm{R}}$ equal $192$.
@@
To investigate this phenomenon, we pre-trained the LLaMA-60M model and collected gradients ${\bm{G}}_{t}$ from different iterations $t$ for examination. Following the setup from GaLore (Zhao et al., 2024a), we computed SVD decompositions and extracted projections ${\bm{P}}_{t}$ with a rank of 128. We evaluated the similarity of the projection matrices by calculating the principal angles between different projections ${\bm{P}}_{t}$ at different steps. Similarly to the observations in Q-Galore (Zhang et al., 2024d), we found that these projections show minimal change during the training period; see Figure 2.
@@
Here, we take the projection matrix corresponding to the $5$-th layer and plot histograms of the cosine of the principal angles between pairs ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations. For comparison, we also include the random projections on the right. As can be seen, the distributions of cosines differ significantly for the ${\bm{P}}_{t}$ and for the random projections. While ${\bm{R}}_{t}$ feature no angles with cosines higher than $0.9,$ the top $57$ cosines for ${\bm{P}}_{t}$ surpass $0.9$, even for gradients $1000$ steps apart.
@@
This leads to the conclusion that although SVD decomposition generally better captures the information contained in the ${\bm{G}}_{t}$, the original GaLore algorithm updates weights only in a small subspace. We hypothesize that training with random projections yields superior results due to the more extensive investigation of the optimizable space during the training process. This finding indicates that to achieve better convergence, it is important to seek out optimization algorithms that explore the entire space during the training process.

### v2

# FRUGAL: Memory-Efficient Optimization by Reducing State Overhead for Scalable Training

@@ Abstract
With the increase in the number of parameters in large language models, the training process increasingly demands larger volumes of GPU memory. A significant portion of this memory is typically consumed by the optimizer state. To overcome this challenge, recent approaches such as low-rank adaptation (LoRA), low-rank gradient projection (GaLore), and blockwise optimization (BAdam) have been proposed. However, in all these algorithms, the effective rank of the weight updates remains low-rank, which can lead to a substantial loss of information from the gradient. This loss can be critically important, especially during the pre-training stage. In this paper, we introduce FRUGAL (Full-Rank Updates with GrAdient spLitting), a new memory-efficient optimization framework. FRUGAL leverages gradient splitting to perform low-dimensional updates using advanced algorithms (such as Adam), while updates along the remaining directions are executed via state-free methods like SGD or signSGD. Our framework can be integrated with various low-rank update selection techniques, including GaLore and BAdam. We provide theoretical convergence guarantees for our framework when using SGDM for low-dimensional updates and SGD for state-free updates. Additionally, our method consistently outperforms concurrent approaches, achieving state-of-the-art results in pre-training and fine-tuning tasks while balancing memory efficiency and performance metrics.
Machine Learning, ICML
@@ 1 Introduction
In recent years, Large Language Models (LLMs) such as GPT (OpenAI, 2023) and LLaMA-3 (Dubey et al., 2024) have demonstrated remarkable performance across various disciplines (Brown, 2020; Yang et al., 2024; Romera-Paredes et al., 2024). However, a critical factor in achieving these results is the size of these models (Hoffmann et al., 2022). Increasing the number of parameters leads to higher computational and memory costs. For example, an 8 billion parameter LLaMA-3 model in 16-bit format requires 32GB just for parameters and gradients. Using the standard Adam optimizer (Kingma, 2014) adds another 32GB for $m$ and $v$ statistics. Moreover, achieving high-quality results often requires 32-bit precision for weights and optimizer states (Zamirai et al., 2020), pushing memory requirements beyond even high-end GPUs like the A100-80GB.
Numerous research projects have been aimed at reducing these significant costs. These approaches include engineering solutions like gradient checkpointing (Chen et al., 2016) and memory offloading (Rajbhandari et al., 2020), which do not change the training trajectory. There are also methods that adjust the training algorithm by decreasing the number of trainable parameters (Frankle & Carbin, 2018; Horváth et al., 2024) or their bit precision (Wortsman et al., 2023), as well as optimizer statistics (Dettmers et al., 2021; Shazeer & Stern, 2018; Zhang et al., 2024c).
Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA (Hu et al., 2021) and Dora (Liu et al., 2024b) reduce memory costs by training a relatively small number of parameters compared to the size of the original model, while the remaining modules are frozen. This approach has proven effective for the task of efficient fine-tuning of pre-trained models. However, PEFT methods have a fundamental limitation: parameter updates always lie in a low-dimensional subspace $L$, which prevents the use of these methods for pre-training (Lialin et al., 2023) and may restrict their capabilities in fine-tuning (Zhang et al., 2024a).
Recent works, such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023) and BAdam (Luo et al., 2024) offer a solution to this problem. These methods enable higher-dimensional full-parameter learning by periodically changing the optimizable low-rank subspace $L$. However, even though these methods result in overall parameter changes that are high-dimensional, the updates in each step remain low-dimensional. The dimensionality of the frozen subspace $\dim M=\dim L^{\perp}$ significantly exceeds $\dim L$. The remaining information contained in the gradient is not utilized for parameter updates. Nevertheless, this information can still be leveraged to train the model.
@@
We present the FRUGAL framework, designed to bridge this gap. Our approach stems from a crucial observation: although memory constraints prevent using optimizers with auxiliary optimizer state — such as Adam (Kingma, 2014) — in the remaining subspace $M$, one still can update $M$ using state-free optimization algorithms like Stochastic Gradient Descent (SGD) or signSGD (Bernstein et al., 2018). This solution allows for high-dimensional updates, which provides additional opportunities to explore the parameter space and improves convergence. We will further refer to the subspaces $L$ and $M$ according to the types of optimizers used for their updates - state-full and state-free.
  • We provide theoretical convergence guarantees for our framework. In the proof, we consider the case with SGDM as the state-full optimizer and SGD as the state-free optimizer, and we show that FRUGAL matches the best-known convergence rate in many scenarios.
  • To verify the practical applicability of FRUGAL, we conduct extensive experiments in popular real-world scenarios [^1] The code is available at https://anonymous.4open.science/r/FRUGAL-D3CA.. In these experiments, we pre-train LLaMA-like models (up to 1B parameters) on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020) and fine-tune RoBERTa (Liu, 2019) on the GLUE benchmark (Wang, 2018). The results show that our method significantly outperforms previous memory-efficient algorithms while using less memory budget.
  • We demonstrate that only the Output layer in transformer-like models requires advanced optimizers like Adam, while other modules (including RMSNorms and Embeddings) can use simpler methods like signSGD without significant performance loss. This opens up new possibilities for memory-efficient training and provides crucial insights into Transformers learning dynamics.
@@
Input: model $f_{\mathbf{\theta}}$ with $p$ parameter sets $\{\theta_{i}\in\mathbb{R}^{d_{i}}\}_{i=1}^{p}$, loss $\mathcal{L}$, gradient projectors $\{P_{k,i}\}_{i=1}^{p},$ number of steps $K$
@@
3:     compute $\ell\leftarrow\mathcal{L}(f_{\theta}(x),y)$Forward}
@@
5:        $g_{\text{full},i}\leftarrow P_{k,i}(g_{i}),\;$Project Grad}
6:        $g_{\text{free, i}}\leftarrow g_{i}-P^{-1}_{k,i}(g_{\text{full},i})$Residual}
7:        $\texttt{s}_{\theta_{i}}\leftarrow[P_{k,i}(P^{-1}_{k-1,i}(s),\;s\in\texttt{s}_{\theta_{i}}]$Project state}
@@ 2 Related work
Memory-efficient full-parameter learning. Recent research has focused on reducing the memory footprint of LLM by decreasing the size of the optimizer states while maintaining their performance. Low-rank adaptation methods, such as LoRA (Hu et al., 2021), inject trainable rank decomposition matrices into linear layers, reducing memory requirements by optimizing only a few learnable adapters. ReLora (Lialin et al., 2023) builds upon this by merging low-rank adaptations into the main model weights during training, increasing the total rank of the update. BAdam (Luo et al., 2024) leverages Block Coordinate Descent for full-parameter training by switching active blocks during fine-tuning. MicroAdam (Modoranu et al., 2024) compresses gradient information before feeding it into the optimizer state, significantly reducing the memory footprint while enabling full parameter learning with error feedback mechanisms. GaLore (Zhao et al., 2024a) maintains full parameter learning by projecting gradients onto a low-rank subspace using SVD decomposition, storing optimizer states in this reduced space. However, while these methods effectively reduce memory overhead, they all perform low-rank updates at each iteration. In contrast, our approach utilizes all available gradient information to perform full-dimensional updates at each optimizer step, offering a novel perspective on memory-efficient optimization for LLM.
However, we note that there also exist several concurrent works — Fira (Chen et al., 2024a), LDAdam (Robert et al., 2024), and Adamem (Vyas et al., 2024) — that also adopt full-dimensional updates for similar goals. See Appendix B for a comparison and detailed discussion.
@@
Other memory-efficient optimization. Several other methods have been proposed to reduce the memory footprint of optimizers. AdaFactor (Shazeer & Stern, 2018) attempts to mimic Adam’s behavior while reducing memory usage through factorization of the variance matrix $v$. Adam-mini (Zhang et al., 2024c) further reduces memory by storing only one value $v$ per block. Dettmers et al. (2021) and Li et al. (2024) decrease memory footprint by quantizing optimizer states to lower-precision representations. Lv et al. (2023) proposed to reduce memory by fusing the backward operation with the optimizer update. Notably, these approaches are orthogonal to our method FRUGAL and can be combined with it for further memory efficiency.
@@
Block Coordinate Descent. Block Coordinate Descent (BCD) is a well-established optimization method with a rich history in mathematical optimization (Ortega & Rheinboldt, 2000; Tseng, 2001; Richtárik & Takáč, 2014, 2015a, 2015b). In recent years, a specific instance of BCD, known as layer-wise learning, has been applied to deep learning. Notable examples include Luo et al. (2024); Pan et al. (2024), which leverage this approach for LLM fine-tuning. To the best of our knowledge, our work presents the first theoretical analysis of an extended BCD framework (Section 5) where the remaining layers are also updated with a different algorithm. This novel approach extends traditional BCD techniques, opening new avenues for full model optimization.
@@
Sign-based methods for training language models. Since its introduction, Adam has become the de facto primary optimization algorithm, demonstrating superior practical results compared to SGD-based algorithms across various deep learning tasks. This difference is particularly noticeable when training Transformers on language tasks. While Zhang et al. (2020) hypothesized that Adam outperforms SGD in this setup due to the heavy-tailed distribution of sampling-induced errors, Kunstner et al. (2023) demonstrated that this superiority persists even in full-batch training. They proposed a new hypothesis suggesting that Adam’s key success factor is related to its similarity to signSGD (Balles & Hennig, 2018; Balles et al., 2020), and both Kunstner et al. (2023) and Zhao et al. (2024b) showed that the signed descent with momentum reduces the performance gap with Adam. In contrast, to the best of our knowledge, we are the first to train the majority of language model parameters using signSGD without momentum, achieving minimal loss in quality. This approach further demonstrates the effectiveness of sign-based methods for LLM training, paving the way for more efficient and scalable optimization strategies.
3 Empirical Analysis and Motivation
@@ 3.1 The importance of exploring the entire space during the training process
In recent work, Zhao et al. (2024a) proposed GaLore, an optimization method based on projecting the gradient matrix ${\bm{G}}$ of each Linear layer222 Since Linear layers contain most parameters and require most memory, we primarly focus on them. onto a low-dimensional subspace. To obtain the projection matrix ${\bm{P}}$, they use the SVD decomposition of ${\bm{G}}_{t}$, which is recomputed with frequency $T$. The vectors or rows of ${\bm{G}}$ are projected onto the first $r$ left or right singular vectors, respectively. This approach has theoretical foundations: the first $r$ singular vectors correspond to the first $r$ singular values and, therefore, should better utilize information from the spectrum of ${\bm{G}}$.
Table 1: Comparison of different projection and state-free subspace optimization strategies on pre-training LLaMA-130M on C4 with AdamW as the state-full algorithm.
Projection Optimizes state- Validation perplexity $\downarrow$
type free subspace 4k 40k 200k
SVD No 39.75 24.38 21.11
Random No 42.31 23.55 20.01
Random Yes 37.26 21.53 18.64
SVD Yes 33.96 21.01 18.35
RandK Yes 36.38 21.25 18.63
Blockwise Yes 37.20 21.42 18.60
AdamW 33.95 20.56 18.13
@@
Given the computational burden of SVD decomposition, a natural question arises about the possibility of employing a random semi-orthogonal projection matrix ${\bm{R}}$ as an alternative to projecting onto the first $r$ singular columns with ${\bm{P}}$. Surprisingly, while the SVD decomposition provides better initial performance, the random projection proves superiority in long-term training, yielding significant improvements. As an illustration, we took the pre-training [^3] See Section 6.1 for a detailed description and discussion. of a 130M model with LLaMA-like architecture on the C4 dataset. The results are presented in the first part of Table 1, where we compare SVD and Random projections.
@@
To investigate this phenomenon, we pre-trained the LLaMA-60M model and collected gradients ${\bm{G}}_{t}$ from different iterations $t$ for examination. We evaluated the similarity of the projection matrices by calculating the principal angles between the projections ${\bm{P}}_{t}$ from different steps. Similarly to the observations in Q-Galore (Zhang et al., 2024d), we found that these projections show minimal change during training; see Figure 2 for details.
@@
Here, we take the projection matrix of k_proj from $5$-th layer and plot histograms of the cosine of the principal angles between pairs ${\bm{P}}_{t}$ and ${\bm{P}}_{t^{\prime}}$ from different iterations. For comparison, we also include the random projections on the right. As can be seen, the distributions of cosines differ significantly for ${\bm{P}}_{t}$ and for ${\bm{R}}_{t}$. While ${\bm{R}}_{t}$ feature no angles with cosines higher than $0.9,$ the top $57$ cosines for ${\bm{P}}_{t}$ surpass $0.9$, even for gradients $1000$ steps apart.
@@
This leads to the conclusion that although the SVD decomposition generally better captures the information contained in ${\bm{G}}_{t}$, the original GaLore algorithm updates the weights only in a small subspace. We hypothesize that training with random projections yields superior results due to the more extensive investigation of the optimizable space during the training process. This finding indicates that to achieve better convergence, it is important to find optimization algorithms that explore the entire space during the training process.